### v1

# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for radiance field rendering, but it typically requires millions of redundant Gaussian primitives, overwhelming memory and rendering budgets. Existing compaction approaches address this by pruning Gaussians based on heuristic importance scores, without global fidelity guarantee. To bridge this gap, we propose a novel optimal transport perspective that casts 3DGS compaction as global Gaussian mixture reduction. Specifically, we first minimize the composite transport divergence over a KD-tree partition to produce a compact geometric representation, and then decouple appearance from geometry by fine-tuning color and opacity attributes with far fewer Gaussian primitives. Experiments on benchmark datasets show that our method (i) yields negligible loss in rendering quality (PSNR, SSIM, LPIPS) compared to vanilla 3DGS with only 10% Gaussians; and (ii) consistently outperforms state-of-the-art 3DGS compaction techniques. Notably, our method is applicable to any stage of vanilla or accelerated 3DGS pipelines, providing an efficient and agnostic pathway to lightweight neural rendering.
Figure 1: Visual comparison. When reducing the number of Gaussians by 90%, our method outperforms other compaction techniques, such as PUP-3DGS, and remains competitive with the original 3DGS.
1 Introduction
Real-time 3D scene reconstruction and rendering dynamically generates photorealistic 3D representations from sensor data (e.g., multi-view images, LiDAR) with minimal latency, enabling critical applications in augmented/virtual reality (AR/VR), autonomous navigation, and immersive media [1, 2]. The current state-of-the-art, 3D Gaussian Splatting (3DGS) [3], iteratively learns 3D anisotropic Gaussian primitives with color and opacity attributes to model these scenes. During rendering, these 3D Gaussians are projected to 2D screens and $\alpha$-blended to achieve real-time photorealistic synthesis.
However, 3DGS faces significant efficiency challenges: its iterative densification process often produces millions of redundant Gaussians for complex scenes [4, 5, 6]. This inefficiency leads to high memory/storage costs and increased per-frame rendering time, limiting deployment on resource-constrained platforms like mobile and AR/VR devices [7].
Figure 2: Comparison of heuristic pruning and our method. The original mixture (left) with $10^{3}$ components is reduced to 5% using either pruning (middle) or our method (right). Our method better preserves the overall structure.
A common strategy to improve efficiency is compaction [8, 9, 10, 7]—reducing the number of Gaussians while preserving rendering fidelity. With fewer primitives, both storage requirements and rendering efficiency (particularly per-frame rendering time) are naturally improved. This approach is viable because 3DGS densification inherently generates redundant primitives [4, 5, 6]. Existing methods achieve compaction via pruning or random subset selection [11, 12, 5, 13, 14, 15, 6, 4, 16]. These strategies, which are based on naïvely discarding information, are simple to implement, but may be ineffective. In particular, they tend to lose critical structural details (as shown in Figure 1) or distort the underlying geometry (as shown in Figure 2). Such losses can degrade rendering quality, particularly in regions with fine-scale features or complex material properties. These limitations motivate our key question:
How to design an efficient compaction method that preserves 3D spatial and structural geometry?
To address this, we frame compaction as an optimization problem, where the goal is to approximate the original 3DGS representation with fewer Gaussians. Our solution leverages a statistical perspective, treating the scene’s geometric structure as a probabilistic model and employing principled reduction techniques to preserve fidelity.
  • (i)
    Geometric compaction via GMR. We first observe that the geometry of a 3DGS representation—defined by the positions, covariances, and opacities of its Gaussian primitives—can be interpreted as a Gaussian mixture model (GMM). Here, the mixture density is a convex combination of individual Gaussian densities, weighted by their opacities. This formulation naturally connects 3DGS compaction to Gaussian Mixture Reduction (GMR), a well-studied problem in statistics where a high-order GMM is approximated by one with fewer components while minimizing a divergence measure.
  • (ii)
    Appearance optimization. While geometry is compacted via GMR, the appearance of the scene—governed by the color attributes of the Gaussians—must also be preserved. To achieve this, we decouple the optimization of geometry (position, covariance) and appearance (color, opacity). After GMR-based compaction, we fine-tune the color and opacity of the reduced set of Gaussians using the standard 3DGS training pipeline. This two-stage strategy ensures that the compacted model maintains both geometric accuracy and photorealistic rendering quality.
For GMR, we minimize the composite transportation divergence [17], which is rooted in optimal transport theory [18, 19] and allows an effective algorithm. We tailor this GMR algorithm (detailed in Section 3.2.1) so that it scales well in scenes like 3DGS with an extensive amount of Gaussians. Crucially, our GMR optimizer does not merely select a subset of existing Gaussians; instead, it creates new primitives that can dynamically adjust their positions and covariances to better approximate the underlying geometry (as shown in Figure 2). Our method is fundamentally algorithm-agnostic: it functions as a plug-and-play module that can enhance both the standard 3DGS pipeline and any of its variants (Section 4), applicable at any stage of training to boost computational efficiency.
To summarize, our contributions are:
  • We open a new pathway to view 3DGS representations as a Gaussian mixture and perform compaction from the perspective of Gaussian mixture reduction via optimal transport.
  • We decouple geometry and appearance optimization in a scalable blockwise optimization scheme.
  • Empirical results demonstrate that our method preserves rendering quality using only 10% Gaussians.
  • We provide a plug-and-play module compatible with any 3DGS variant at any training stage.
2 Related Works
Compaction Techniques in 3DGS 3DGS compaction seeks to minimize the Gaussian count while preserving image quality. Existing work falls into two operations: densification (where to add) and pruning (what to drop). Most current approaches rely on per-Gaussian heuristic scores.
For densification, Taming-3DGS [10] ranks candidate Gaussians by combining gradient, pixel coverage, per-view saliency, and core attributes including opacity, depth, and scale; Color-cued GS [20] considers the view-independent spherical harmonics coefficient gradient to better capture color cues; and GaussianPro [9] guides growth using depth and normal maps. In the pruning phase, LightGaussian [11], Mini-Splatting [12], RadSplat [5], and AtomGS [13] compute an importance score from each Gaussian’s accumulated ray contribution, typically a mix of volume, opacity, transmittance, and hit count, and discard the lowest-ranked Gaussians. Gradient-aware variants prune by per-Gaussian gradients (Trimming-the-Fat) [14] or the second-order sensitivity score derived from the Hessian matrix (PUP-3DGS) [15]. The score can also be trained via a learnable mask, as in LP-3DGS [6], Compact3DGS [4], and HAC [16]. Moreover, multi-view consistency criteria discard Gaussians unseen by keyframes [21] or visible only in real but not virtual views [22]. For a comprehensive survey of 3DGS compression and compaction, we refer readers to [8, 23].
Despite their success, most existing strategies evaluate each Gaussian independently, leaving open the question of whether the retained set is truly the best global surrogate. Our work addresses this gap from a probabilistic perspective via Gaussian mixture reduction.
3DGS from Probabilistic Distribution Point of View Kheradmand et al. [24] formulate 3DGS as a Markov chain Monte Carlo process and use stochastic gradient Langevin dynamics to migrate dropped Gaussians onto retained ones, partially recycling lost information. However, their update remains pairwise and lacks convergence guarantees under a principled divergence. Moreover, this scheme is coupled with the original 3DGS pipeline, limiting its generalizability to other variants.
3 Method: Gaussian Herding across Pens
Figure 3: An illustration of the proposed GHAP-3DGS approach. The process begins with full-resolution 3DGS training to obtain initial geometric and appearance features. These Gaussians are then spatially partitioned using a KD-tree and grouped into blocks—analogous to sheep pens. We then perform blockwise Gaussian Model Reduction (GMR) to generate a small number of representative Gaussians within each block. This step is analogous to the popular kernel herding method  [25]. Finally, a lightweight appearance refinement step further optimizes the appearance feature of the reduced set. This multi-stage pipeline progressively guides the Gaussians in each block—analogous to herding across pens—toward a compact and high-fidelity representation.
We present a principled approach for compacting 3DGS representations through Gaussian Mixture Reduction (GMR). Our key insight is that The 3DGS densification strategy, while effective for view synthesis, produces redundant Gaussian components that can be consolidated without sacrificing visual quality. We achieve this by interpreting the 3D scene surface as an unnormalized probability density function (PDF), where the trained 3DGS representation serves as a Gaussian mixture approximation of this PDF. Thus, our problem reduces to finding a Gaussian mixture with fewer components to approximate the original, redundant one. Our procedure separately optimizing geometric (position $\mu$, covariance $\Sigma$) and appearance (opacity $\alpha$, color $c$) attributes to ensure rendering fidelity. The key idea of our method is illustrated in Figure 3.
3.1 Probabilistic Scene Representation
Let $\phi(x;\mu,\Sigma)=|2\pi\Sigma|^{-1}\exp(-(x-\mu)^{\top}\Sigma^{-1}(x-\mu))$ be the PDF of a Gaussian distribution with mean $\mu$ and covariance $\Sigma$. A Gaussian mixture with $n$ components is a distribution with density:
$\phi_{n}(x)=\sum_{i=1}^{n}\alpha_{i}\phi(x;\mu_{i},\Sigma_{i}),$
[…]
$\phi_{n}(x)\approx f(x),\quad\forall x\in{\mathcal{X}},$
where ${\mathcal{X}}$ is the 3D scene volume. Therefore, the geometry of the 3D scene can be effectively represented by a Gaussian mixture. Then, each of these Gaussian primitive is associated with its own color $c_{i}$. Both the geometry and appearance attributes are important for high quality rendering.
3.2 Compaction via Optimal Transport
Motivated by the observation that many 3DGS algorithms [4, 5, 6] produce a significant number of redundant Gaussians during training, we improve rendering efficiency through compaction—reducing the number of Gaussian primitives to achieve lower memory usage and faster rendering while preserving visual fidelity. The process consists of two key phases:
  1. 1.
    Geometric Compaction via GMR: Leveraging our probabilistic interpretation, we formulate compaction as Gaussian Mixture Reduction (GMR) [26], approximating the original Gaussian mixture with redundant components with fewer components. This yields a compacted geometric representation: $\bar{\phi}_{m}(x)=\sum_{j=1}^{m}\bar{\alpha}_{j}\phi(x;\bar{\mu}_{j},\bar{\Sigma}_{j})$, where $m\ll n$. This step modifies only the Gaussian positions ($\bar{\mu}_{j}$) and covariances ($\bar{\Sigma}_{j}$), leaving appearance attributes unchanged.
  2. 2.
    Appearance Optimization: The reduced Gaussians are initialized with appearance attributes (colors, opacities) and fine-tuned for optimal rendering performance. This step optimizes appearance only, maintaining geometric consistency.
Our approach decouples geometry and appearance optimization while using standard 3DGS training to preserve quality. Our training pipeline is visualized in Figure 3 and we describe the details below.
3.2.1 Geometric Compaction via GMR
Following Zhang et al. [17], we formulate compaction as minimizing the composite transportation divergence (CTD) between two Gaussian mixtures:
Definition 1 (Composite transportation divergence).
Let $c(\cdot,\cdot)$ be a divergence between two Gaussians. The composite transportation divergence (CTD) between two Gaussian mixtures $\phi_{n}(x)=\sum_{i=1}^{n}\alpha_{i}\phi(x;\mu_{i},\Sigma_{i})$ and $\phi^{\prime}_{m}(x)=\sum_{j=1}^{m}\alpha^{\prime}_{j}\phi(x;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$ with cost function $c(\cdot,\cdot)$ is
${{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m})=\inf\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})):\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\right\}.$
The CTD generalizes optimal transport [18] to mixtures, treating each component as a discrete distribution in Gaussian space. The cost function measures the cost of moving one unit of Gaussian from one location to another, and $\pi_{ij}$ measures the corresponding amount of mass that is being moved. The total cost is proportional to the cost and the mass, and the divergence is the smallest transportation cost to move the original mixture to the target mixture. The reduced mixture becomes
$\{\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j}\}=\operatorname*{arg\,min}_{\{\alpha^{\prime}_{j},\mu^{\prime}_{j},\Sigma^{\prime}_{j}\}}{{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m}).$ (1)
With this formulation, the Gaussian mixture after compaction has optimal guarantee. The solution also guides the choice of $m$ to balance compactness and fidelity.
Algorithm 1 GMR via $k$-means Clustering
1:Initialize $\{\bar{\mu}_{j}^{(0)},\bar{\Sigma}_{j}^{(0)}\}_{j=1}^{m}$
2:for t=1,…, do
3:     Assignment Step:
4:     for $i=1$ to $n$ do
5:         Assign component $i$ to cluster ${\mathcal{C}}_{j}$ that minimizes $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j}^{(t-1)},\bar{\Sigma}_{j}^{(t-1)}))$
6:     end for
7:     Update Step:
8:     for $j=1$ to $m$ do
9:         Compute new cluster center: $\bar{\mu}_{j}^{(t)},\bar{\Sigma}_{j}^{(t)}=\operatorname*{arg\,min}\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}c(\phi(\cdot,\mu_{i},\Sigma_{i}),\phi(\cdot,\mu,\Sigma))$
10:     end for
11:     if no change in assignments then
12:         for $j=1$ to $m$ do
13:              Compute weight: $\bar{\alpha}_{j}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}$
14:         end for
15:         break
16:     end if
17:end for
As shown in Zhang et al. [17], (1) can be solved using the effective iterative algorithm in Algorithm 1. The algorithm reduces to a $k$-means variant in Gaussian space: 1) The assignment step follows the same principle as traditional $k$-means, but replaces the $L^{2}$ distance between vectors with a cost function $c(\cdot,\cdot)$ between Gaussian distributions. 2) The update step generalizes the cluster center computation: In traditional $k$-means, centers are updated as arithmetic averages (barycenters w.r.t. $L^{2}$ distance) of vectors in each cluster. In this algorithm, centers become barycenters of Gaussians in each cluster, minimized w.r.t. cost function $c(\cdot,\cdot)$. Thus, standard $k$-means emerges as a special case when using $L^{2}$ distance on vectorized Gaussian parameters.
While the standard GMR algorithm provides optimal theoretical guarantees, its direct application to 3DGS compaction proves computationally prohibitive. Although the algorithm must converge in finite steps [17] [^2] In particular, we find that in our experiment, the algorithm converges in only around 6 iterations., the assignment step involves $nm$ evaluations of the cost function per iteration. In typical 3D scenes, the number of Gaussians scales as $n=\Omega(10^{5})$ [^3] The $a_{n}=\Omega(b_{n})$ ($O(b_{n})$) if there exist $C\geq 0$ such that $a_{n}\geq Cb_{n}$ ($a_{n}\leq Cb_{n}$)., and even after 95% reduction, each iteration would still require at least $10^{8}$ operations and memory storage. For the update step, the computational cost for the cluster center depends on the pre-specified cost function $c(\cdot,\cdot)$. The KL divergence considered in Zhang et al. [17] suffers from its key limitations: it requires computing inverse covariance matrices, which introduces (1) numerical instability due to small eigenvalues of covariance matrices, and (2) significant overhead despite operating in ${\mathbb{R}}^{3}$ (with $O(\rho s^{2}\log n)$ inversions needed). To overcome this challenge, we introduce two key optimizations designed for 3DGS:
  • Blockwise GMR via KD-Tree: To improve computational and memory efficiency during training, we partition the scene into spatially localized blocks and perform GMR independently within each block. As demonstrated in Remark 1, this blockwise approach yields significant computational savings. While both KD-trees [27] and Octrees [28] are effective for spatial partitioning in 3D space, we employ a KD-tree for two key advantages. First, it produces more balanced partitions across regions. Second, it avoids unnecessary subdivisions in sparse regions that would waste computational resources.
    Our KD-tree is constructed solely from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ (justified by the observed small eigenvalues of covariance matrices). Each split uses the median coordinate value, creating $2^{d}$ blocks at depth $d$. We set $d=\lfloor\log_{2}(n/s)\rfloor$ to ensure blocks contain at most $s$ Gaussians with $s\ll n$, then reduce each block to $m=\rho s$ components ($\rho$ = retention ratio).
    Remark 1 (Computational cost comparison).
    Our blockwise approach reduces the per-iteration computational cost from $O(\rho n^{2})$ to $O(\rho s^{2})$ per block. [^4] Each block reduces from $s$ Gaussian components to $\rho s$. With $2^{\text{depth}}=O(\log n)$ blocks in total, the overall complexity becomes $O(\rho s^{2}\log n)$. For typical values of $n=10^{5}$, $s=10^{3}$ and $\rho=0.05$, this reduces the cost from $10^{8}$ to approximately $10^{5}$ operations–a substantial improvement. The savings become even more pronounced for larger $n$. Furthermore, the reduction steps can be executed in parallel across blocks, offering additional computational speedup.
  • Efficient Cost Function: We propose a novel cost function that is both computationally efficient and maintains approximation quality, overcoming limitations of KL divergence used in [17]. We propose to use the following divergence:
    $c(\phi(\cdot;\mu,\Sigma),\phi(\cdot;\mu^{\prime},\Sigma^{\prime}))=\|\mu-\mu^{\prime}\|_{2}^{2}+\|\Sigma-\Sigma^{\prime}\|_{F}^{2},$ (2)
    which has three key advantages: First, it preserves distributional similarity as Gaussians are uniquely determined by their mean and covariance. Second, the assignment step requires only efficient vector/matrix norm computations. Third, the update step simplifies to weighted averages:
    $\bar{\mu}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\mu_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i},\quad\quad\quad\bar{\Sigma}_{j}^{(t)}=\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}\Sigma_{i}/\sum_{i\in{\mathcal{C}}_{j}}\alpha_{i}.$
3.2.2 Appearance Optimization
Algorithm 2 GHAP: 3DGS Compaction via Blockwise GMR
1:Input: Trained 3DGS model for $T$ steps to obtain $\{(\alpha_{i},\mu_{i},\Sigma_{i},c_{i})\}_{i=1}^{n}$, retention ratio $\rho$
2:Output: Compacted $\{(\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j},\bar{c}_{j})\}_{j=1}^{m}$
3:Stage 1: Geometric Compaction
4:1. Build KD-tree from Gaussian centers $\{\mu_{i}\}_{i=1}^{n}$ with depth $d=\lfloor\log_{2}(n/s)\rfloor$
5:2. For each leaf block ${\mathcal{B}}_{k}$:
6: Run Algorithm 1 to reduce to $\rho s$ Gaussians
7:Stage 2: Appearance Optimization
8:1. Initialize appearance for each $\bar{\phi}_{j}$:
9:$\bar{c}_{j}\leftarrow c_{i^{*}}$ and $\bar{\alpha}_{j}\leftarrow\alpha_{i^{*}}$ where $i^{*}=\operatorname*{arg\,min}_{i\in[n]}\|\mu_{i}-\bar{\mu}_{j}\|_{2}$
10:2. Fine-tune $\{\bar{\alpha}_{j},\bar{c}_{j}\}$ using standard 3DGS rendering pipeline for $T$ steps
Following geometric compaction, we initialize the appearance attributes (opacity and colors) of the compacted Gaussian primitives. For each primitive in the reduced mixture, we assign the appearance parameters from its closest counterpart in the original Gaussian mixture. Using these initial values, we then optimize the appearance attributes through backpropagation following the standard 3DGS training pipeline. We optimize the opacity instead of directly using the opacity from the GMR algorithm since these output weights not necessarily satisfy the constraint on the opacity that it must be within 00 and $1$. Finetuning the the opacity can achieve better visualization performance.
3.3 Training Details with GHAP Algorithm
Integrating these components, we present our complete training pipeline in Algorithm 2, called Gaussian Herding Across Pens (GHAP). The process begins with standard 3DGS optimization for $T$ steps, followed by blockwise GMR. We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization. This procedure can be iteratively applied throughout training as needed.
4 Experiments
4.1 Experimental Setup
Datasets
For a comprehensive evaluation of our GHAP algorithm, we use three real-world datasets: Tanks and Temples [29], Mip-NeRF 360 [30], and Deep Blending [31], which cover varying levels of detail, lighting conditions, and scene complexities. For each dataset, we adopt the same scenes as in [8].
  • Tanks and Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
  • Mip-NeRF 360: We test on a mix of indoor and outdoor scenes, including “Bicycle”, “Bonsai”, “Counter”, “Flowers”, “Garden”, “Kitchen”, “Room”, “Stump”, and “Treehill”, all with centered viewpoints.
  • Deep Blending: We include two indoor scenes, “Dr. Johnson” and “Playroom”, where the viewpoint is directed outward.
Baselines
To evaluate the effectiveness of our proposed method, we compare it against several strong baselines in the domain of 3DGS compaction: Mini-Splatting-D [12], AtomGS [13], LightGaussian [11], and PUP-3DGS [15]. We set $s=3\times 10^{3}$. The sensitivity analysis of $s$ is given in Appendix. The initialization of the iterative algorithm in (1) randomly selects $m$ Gaussians per block.
Evaluation Metrics We assess 3DGS compaction using standard metrics for rendering quality and efficiency. For quality, we report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model. Higher PSNR/SSIM and lower LPIPS indicate better quality. For efficiency, we measure FPS across retention ratios, where lower retention yielding higher FPS implies greater rendering efficiency.
Device
All experiments were conducted on a server with 256 GB RAM and a 96-core Intel® Xeon® Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
4.2 Quantitative Results
We first demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants. Second, we show that our approach outperforms other SOTA compaction techniques. We present the results first and summarize our findings at the end of the section.
As a Plug-In Compaction Method
Our method can be used as a plug-in compaction method in various 3DGS training algorithms. This demonstrates the broad applicability of our proposed method. To verify this, we apply our compaction method within various 3DGS pipelines. Specifically, we consider three representative variants: 3DGS-30K, AtomGS, and Mini-Splatting-D. Each of them employs a distinct densification strategy, and our method can be directly embedded into the pipeline without extensive engineering effort.
For each variant, we evaluate performance under two retention ratios (10% and 20%) and report results using three standard metrics: PSNR, SSIM, and LPIPS. Table 1 summarizes our experimental results.
Table 1: Quantitative evaluation of our GHAP method with respect to compaction performance applied to different 3DGS variants under varying retention ratios (10% and 20%).
Method $\rho$ Tanks&Temples MipNeRF-360 Deep Blendering
[…]
3DGS-30k 10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
Mini-Splatting-D 10% 0.835 23.232 0.198 0.802 27.090 0.250 0.909 30.042 0.254
20% 0.855 23.403 0.171 0.821 27.310 0.214 0.912 30.170 0.238
100% 0.848 23.338 0.140 0.832 27.486 0.176 0.907 29.980 0.211
AtomGS 10% 0.793 22.988 0.274 0.764 26.535 0.307 0.899 29.347 0.282
20% 0.812 23.282 0.240 0.788 27.025 0.269 0.902 29.347 0.268
100% 0.814 23.289 0.235 0.796 27.135 0.251 0.902 29.314 0.267
Comparison with SOTA Compaction Methods
We compare our method with two recent compaction baselines: PUP-3DGS and LightGaussian. To ensure a fair evaluation, all methods are built upon the same backbone model, 3DGS-30K. Compaction is performed at the same stage, when the number of Gaussians reaches 15K, followed by 15K fine-tuning iterations. During fine-tuning, we use identical optimization settings across all methods, including learning rate and other hyperparameters. Two compaction ratios are considered: 10% and 20%. A detailed comparison of rendering quality across different compaction methods and retention ratios is presented in Table 2.
Table 2: We compare our GHAP method against LightGaussian and PUP-3DGS, modifying their original workflows to align with our compaction pipeline.
Method-Compact Rate Tandt&Temples MipNeRF-360 Deep Blendering
[…]
3DGS-30k-100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
GHAP(Ours)-10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
Light-GS-10% 0.791 22.560 0.248 0.779 26.585 0.261 0.861 27.136 0.326
Pup-GS-10% 0.767 21.519 0.28 0.753 25.332 0.309 0.895 29.153 0.274
GHAP(Ours)-20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
Light-GS-20% 0.791 22.559 0.248 0.779 26.588 0.261 0.861 27.142 0.326
Pup-GS-20% 0.809 22.603 0.228 0.790 26.671 0.257 0.905 29.719 0.248
Figure 4: Visualization before and after compaction. Various scenes before and after compaction under different 3DGS algorithms. Our GHAP method yields negligible loss in rendering quality.
Takeaways
Based on these experiments, we draw four key observations:
  • Superior fidelity retention. Our method substantially outperforms both LightGaussian and PUP-3DGS in SSIM and PSNR. Surprisingly, our method can even achieve better performance than the original method, this is demonstrated in Table 1 on the Deep Blending scene with the Mini-Splatting-D backbone and at a 10% retention ratio. Likewise, in Table 2, we consistently surpass LightGaussian and PUP-3DGS in most datasets and metrics.
  • Efficiency gains. At a 10% Gaussian primitives retention rate, our compaction framework yields a 140.3% increase in rendering speed (FPS), demonstrating that significant model slimming can be achieved without sacrificing and even often improving performance (see ablation study).
  • Cross-dataset consistency. Performance remains stable across datasets. In outdoor scenes (Tanks and Temples), SSIM value fidelity drops ratio (relative drop in rendering metrics after compaction) less than 4.1%. Meanwhile, for the indoor scene, the PSNR value retains 99.4% of the original value.
  • LPIPS sensitivity. We observe that our GMR compaction method incurs a relatively larger degradation in LPIPS compared to SSIM and PSNR, although the reduction remains within acceptable bounds. This trend aligns with prior findings [11, 15], as LPIPS is more sensitive to structural and textural distortions. Since compaction inherently reduces model capacity, fine-grained texture reconstruction becomes more challenging, which explains the behavior observed in our qualitative and quantitative analyses
Qualitative Visualization
We visualize two scenes for each baseline 3DGS algorithm before and after compaction in Figure 4. Across most scenes, our compaction method successfully preserves rendering quality using only 10% of the Gaussian primitives. By design, our reduced Gaussians approximate the original Gaussians from the 3DGS algorithm. If the original model lacks good visual quality due to lack to Gaussian primitives in certain regions, our method cannot perform well—as demonstrated in the ‘Bicycle’ scene, where the poor performance of the Atom-GS baseline prevents our approach from maintaining high fidelity after compaction. Interestingly, in some cases, our post-compaction results surpass the pre-compaction quality. A representative example is the ‘Kitchen’ scene based on Mini-Splatting-D. We conjecture this occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an oversaturated set of Gaussians that introduce visual artifacts (e.g., unnatural shadows in the lower-left region). Our method mitigates this issue by globally regularizing the Gaussian distribution: improving expressiveness, reducing unnecessary density, and preserving the underlying 3D structure.
4.3 Ablation Studies
GMR versus Random Subsampling We evaluate the contribution of each stage in our pipeline—geometric compaction and appearance optimization—on the Tanks and Temples dataset. For comparison, we use random subsampling at the same compaction ratio. This allows us to assess the effectiveness of our design choices at each stage. The experiments show that our compaction method is significantly more effective than random subsampling,
Table 3: Ablation study. Mean PSNR, SSIM, and LPIPS on the Tanks and Temples dataset at each stage of GHAP pipeline, with random subsampling as a control. Each operation is applied cumulatively to all subsequent stages.
Method Tanks&Temples
[…]
3DGS-15K 0.839 23.084 0.194
+ 10% GHAP Compaction 0.502 14.015 0.483
+ 10% Random Compaction 0.333 9.573 0.555
+15K Fine-tuning 0.818 23.312 0.242
+15K Fine-tuning 0.712 21.312 0.282
producing a more meaningful scene representation. Appearance optimization further leads to substantial quantitative improvements. Together, these results validate the necessity and effectiveness of both stages in our proposed pipeline.
Figure 5: GHAP efficiency-quality trade-off. SSIM and FPS (top) and LPIPS and PSNR (bottom) across different retention ratios $\rho$.
Rendering Metrics under GHAP
To analyze the efficiency-quality trade-off of the proposed GHAP approach, we vary the retention ratio $\rho$ from $10\%$ to $100\%$. A lower retention ratio enables more aggressive compaction, resulting in faster rendering speed but potentially compromising rendering quality. Our results demonstrate that decreasing the retention ratio yields 2×2$\times$2 × faster rendering with only marginal quality degradation. For instance, as $\rho$ decreases from $100\%$ to $10\%$: LPIPS increases modestly from 0.14 to 0.21; SSIM declines slightly from 0.885 to 0.860; PSNR remains stable between 25.0 dB and 25.5 dB. These findings indicate that retaining even a modest number of Gaussians effectively preserves visual quality, achieving an excellent performance-fidelity trade-off.
5 Conclusion
We present a novel optimal transport-based global Gaussian reduction approach for 3D Gaussian Splatting, addressing the inefficiencies of existing Gaussian pruning and compaction methods. By formulating the problem as Gaussian mixture reduction, our method minimizes composite transport divergence to derive a compact geometric representation, followed by appearance fine-tuning with significantly fewer primitives. It provides a theoretically sound foundation for efficient neural rendering. Extensive experiments demonstrate that our approach achieves rendering quality nearly indistinguishable from vanilla 3DGS while retaining only  10% of the Gaussians, outperforming existing compaction techniques. Numerical results show our framework is versatile, compatible with various 3DGS pipelines. This work advances the practical deployment of 3DGS by striking an optimal balance between computational efficiency and visual fidelity. Future work includes extending our method to dynamic 3DGS for real-time temporal scenes and developing adaptive reduction strategies for task-specific needs.
Acknowledgments and Disclosure of Funding
Qiong Zhang is supported by the National Natural Science Foundation of China Grant 12301391.
Appendix
Appendix A More Details about the Blockwise GMR Compaction
A.1 Introduction to Optimal Transport (OT)
The Optimal Transport (OT) problem dates back to 1781, when French mathematician Gaspard Monge (1746–1818) first formulated it as finding the minimal-cost way to move sand into a hole. This “transport plan” minimizes the transportation cost, hence the term “optimal transport”. The original formulation is known as Monge’s problem. In 1947, Russian economist Leonid Kantorovich (1912–1986) relaxed Monge’s formulation, leading to the so-called Monge-Kantorovich problem.
Specifically, let $P$ and $Q$ be two distributions over a metric space ${\mathcal{X}}$, let $c:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}_{+}$ be a cost function, and let the coupling
$\Pi(P,Q)=\left\{\pi(x,y):\int\pi(dx,\cdot)=Q(\cdot),\int\pi(\cdot,dy)=P(\cdot)\right\}$
be the set of joint distributions with marginals $P$ and $Q$. For any cost function $c$, the total transportation cost induced by a plan $\pi\in\Pi(P,Q)$ is defined as
${\mathcal{I}}_{c}(\pi)=\int_{\mathcal{X}\times\mathcal{X}}c(x,y)\,\pi(dx,dy).$
Here, $\pi(x,y)$ indicates how much “mass” is transported from $x$ to $y$. The first constraint, $\int\pi(x,dy)=P(x)$, ensures that the mass at location $x$ is spread over ${\mathcal{X}}$, while the second constraint, $\int\pi(dx,y)=Q(y)$, ensures that the destination at $y$ receives the required mass. The OT problem seeks the optimal plan
$\pi^{*}=\operatorname*{arg\,min}\left\{{\mathcal{I}}_{c}(\pi):\pi\in\Pi(P,Q)\right\},$
minimizing the transportation cost. The corresponding minimal cost,
$\mathcal{T}_{c}(P,Q)={\mathcal{I}}_{c}(\pi^{*}),$
induces a divergence between $P$ and $Q$. This OT divergence provides a principled way to measure distributional similarity, enabling applications in density matching, generative modeling, and dimensionality reduction.
The optimal plan $\pi^{*}$ and OT divergence typically lack closed-form solutions. Numerical algorithms [18] are used to compute the OT between discrete measures. When $P$ and $Q$ are discrete, the OT divergence reduces to:
Example 1 (OT Divergence Between Discrete Distributions).
For $P=\sum_{i=1}^{n}u_{n}\delta_{x_{n}}$ and $Q=\sum_{j=1}^{m}v_{m}\delta_{y_{m}}$, the OT divergence is
${\mathcal{T}}_{c}(P,Q)=\min\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(x_{i},y_{j}):\sum_{i=1}^{n}\pi_{ij}=v_{j},\sum_{j=1}^{m}\pi_{ij}=u_{i}\right\}.$ (3)
Exact solutions can be found via linear programming, while approximations are obtained using algorithms like Sinkhorn.
A.2 Relationship Between Composite Transportation Divergence and OT
The Composite Transportation Divergence (CTD) [32, 33] extends OT to Gaussian mixtures. For two mixtures $\phi_{n}=\sum_{i=1}^{n}\alpha_{i}\phi(\cdot;\mu_{i},\Sigma_{i})$ and $\phi^{\prime}_{m}=\sum_{j=1}^{m}\alpha^{\prime}_{j}\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$, the CTD is:
${{\mathcal{T}}}_{c}(\phi_{n},\phi^{\prime}_{m})=\inf\left\{\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})):\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\right\},$
Comparing this with the discrete OT divergence (3), the CTD treats Gaussian mixtures as discrete distributions over the space of Gaussians, defining the divergence as the OT between them.
To illustrate, consider $n$ warehouses and $m$ factories in the space of Gaussian distributions ${\mathcal{F}}$. The $i$th warehouse, at $\phi(\cdot;\mu_{i},\Sigma_{i})$, holds $\alpha_{i}$ units of material, while the $j$th factory, at $\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j})$, requires $\alpha^{\prime}_{j}$ units. The cost to transport material from $i$ to $j$ is $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j}))$, and $\pi_{ij}\geq 0$ denotes the transported amount. The total cost under plan $\pi$ is $\sum_{i,j}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\mu^{\prime}_{j},\Sigma^{\prime}_{j}))$. The coupling set $\Pi(\alpha,\alpha^{\prime})=\{\pi_{ij}:\sum_{j=1}^{m}\pi_{ij}=\alpha_{i},\sum_{i=1}^{n}\pi_{ij}=\alpha^{\prime}_{j}\}$ ensures: (a) correct material removal from warehouses, and (b) correct delivery to factories. The OT problem seeks the plan $\pi^{*}\in\Pi(\alpha,\alpha^{\prime})$ minimizing the total cost. The minimal cost corresponds to the CTD between the mixtures, quantifying the optimal transport cost between them.
Our compaction method leverages this interpretation, approximating a mixture with fewer Gaussians by minimizing their CTD-based dissimilarity.
A.3 Algorithm Intuition
At first glance, the optimization problem in (1) is bilevel: the OT plan must be found for any candidate $\{\bar{\alpha}_{j},\bar{\mu}_{j},\bar{\Sigma}_{j}\}_{j=1}^{m}$, and the objective must be minimized. However, as shown by Zhang et al. [17], this simplifies with a clear interpretation. The column-wise marginal constraints on $\pi$ are redundant, and each Gaussian primitive $(\mu_{i},\Sigma_{i})$ transports its full mass to its “closest” counterpart. The optimal plan thus corresponds to clustering $\{(\mu_{i},\Sigma_{i})\}_{i=1}^{n}$ into $m$ clusters $\{{\mathcal{C}}_{j}\}_{j=1}^{m}$, with the original Gaussians forming cluster “barycenters”. Mathematically, the simplified program is:
$\min\left\{\sum_{j=1}^{m}\sum_{i\in{\mathcal{C}}_{j}}\pi_{ij}c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j},\bar{\Sigma}_{j}))\colon[n]={\mathcal{C}}_{1}\sqcup\cdots\sqcup{\mathcal{C}}_{m}\right\},$
where $\sqcup$ denotes disjoint union. The optimal plan has a closed form:
$\pi_{ij}=\begin{cases}\alpha_{i}&\text{if }j=\operatorname*{arg\,min}_{k}c\left(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{k},\bar{\Sigma}_{k})\right),\\ 0&\text{otherwise}.\end{cases}$
With this, $\pi$ and the reduced mixture parameters can be updated alternately, formalized in the following $k$-means-like algorithm:
  • Assignment Step: Each Gaussian $\phi(\cdot;\mu_{i},\Sigma_{i})$ is assigned to cluster $\mathcal{C}_{j}$ by minimizing $c(\phi(\cdot;\mu_{i},\Sigma_{i}),\phi(\cdot;\bar{\mu}_{j},\bar{\Sigma}_{j}))$, analogous to $k$-means’ nearest-centroid assignment.
  • Update Step: For the cost function
    $c(\phi,\phi^{\prime})=\|\mu-\bar{\mu}\|_{2}^{2}+\|\Sigma-\bar{\Sigma}\|_{F}^{2},$
    the parameters $(\bar{\mu}_{j},\bar{\Sigma}_{j})$ are updated as weighted averages:
    $\bar{\mu}_{j}=\frac{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}\mu_{i}}{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}},\quad\bar{\Sigma}_{j}=\frac{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}\Sigma_{i}}{\sum_{i\in\mathcal{C}_{j}}\alpha_{i}}.$
    This mirrors $k$-means’ centroid update.
A.4 Algorithm Convergence
The following theorem guarantees convergence of the CTD sequence generated by Algorithm 1. In the worst case, the algorithm requires exponentially many steps, but it typically converges in about 5 iterations. A full proof is given in Zhang et al. [17].
Theorem 1 (Convergence of the Algorithm).
Suppose $c(\cdot,\cdot)$ is continuous, and for any $\Delta>0$ and $\phi^{\star}$, the set $\{\phi:c(\phi^{*},\phi)\leq\Delta\}$ is compact under some Gaussian space metric. Let $\{\bar{\phi}_{m}^{(t)}\}$ be the sequence generated by the update step with initial $\{\bar{\alpha}_{j}^{(0)},\bar{\mu}_{j}^{(0)},\bar{\Sigma}_{j}^{(0)}\}$, and let $\mathcal{T}_{c}^{(t+1)}$ be the CTD at iteration $t$. Then:
  1. 1.
    There exists $T$ and $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ such that for all $t\geq T$, $\{\bar{\alpha}_{j}^{(t)},\bar{\mu}_{j}^{(t)},\bar{\Sigma}_{j}^{(t)}\}=\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ and $\mathcal{T}_{c}^{(t+1)}=\mathcal{T}_{c}^{(*)}$, where $\mathcal{T}_{c}^{(*)}$ is the CTD between the original mixture and $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$.
  2. 2.
    The limit point $\{\bar{\alpha}_{j}^{*},\bar{\mu}_{j}^{*},\bar{\Sigma}_{j}^{*}\}$ is a local minimum of ${\mathcal{T}}_{c}$.
  3. 3.
    An MM-based exhaustive algorithm with $O(m^{n})$ complexity solves (1).
Appendix B More Background and Involved Techniques in Paper
B.1 Gaussian splatting and efficient variants
Due to its advantages in rendering speed and visual fidelity, 3DGS has been adopted in a broad range of applications, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37]. Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40]. However, the standard 3DGS suffers from significant storage, memory, and rendering costs when applied to large-scale or complex scenes. This limits its deployment on resource-constrained platforms and in broader areas.
To improve storage and rendering efficiency while maintaining visual quality, recent research has focused on two complementary directions: compression and compaction [8]. Compression methods shrink file size by exploiting redundancy in Gaussian attributes (e.g., using vector quantization to cluster similar parameters) [11, 41] or structuring Gaussians into more compact forms (e.g., grids, anchors, and shared codes) [42, 43]. On the other hand, compaction aims to reduce the number of Gaussians required for accurate rendering by optimizing their spatial distribution and pruning redundant primitives [12, 6, 5]. Hybrid pipelines such as LightGaussian [11] and Octree-GS [7] combine compression and compaction together to deliver highly compact and lightweight neural renders.
B.2 Detailed 3DGS pipline
The advantages of 3DGS in rendering speed and image fidelity have made it applicable to a wide range of tasks, including human reconstruction, AI-generated content, autonomous driving, and beyond [34, 35, 36, 37]. Extensions to dynamic 3DGS, editable 3DGS, and surface representation have further broadened its utility [38, 39, 40]. 3D Gaussian Splatting (3DGS) represents a scene as a set of anisotropic 3D Gaussian primitives, each parameterized by its spatial location, shape, opacity, and radiance. The pipeline consists of the following key steps:
  1. 1.
    Initialization. From Structure-from-Motion (SfM), obtain calibrated camera poses and a sparse point cloud. Each point is initialized as a 3D Gaussian with an opacity $\alpha_{i}$:
    $\phi(x;\mu_{i},\Sigma_{i})=\left|2\pi\Sigma\right|^{-1}\exp\left(-\frac{1}{2}\left(x-\mu_{i}\right)^{T}\Sigma_{i}^{-1}\left(x-\mu_{i}\right)\right),$
    where $\mu_{i}\in\mathbb{R}^{3}$ is the position (mean), $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix (anisotropic shape), and $\alpha_{i}\in[0,1)$.
  2. 2.
    Projection & Rasterization. Each 3D Gaussian is projected to 2D using the camera model,
    $\Sigma_{i}^{\prime}=J{W}\Sigma_{i}J{W}^{T},$
    where $W$ is a veiw transformation matrix and $J$ is the Jacobian of the projective transform. Rasterization is done using a differentiable splatting approach which makes optimization possible.
  3. 3.
    Image Formation (Alpha Blending). The pixel color is computed via volumetric blending:
    $C=\sum_{i=1}^{N}T_{i}\cdot\alpha_{i}\cdot c_{i}\text{ with }T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$
    where $c_{i}$ is the SH-predicted color of the $i$-th Gaussian in front-to-back order.
  4. 4.
    Optimization. The Gaussian parameters $\left\{\mu_{i},\Sigma_{i},\alpha_{i},c_{i}\right\}$ are optimized to minimize a photometric loss:
    $\mathcal{L}=\left(1-\lambda\right)\left\|\widehat{\bm{C}}-\bm{C}^{\star}\right\|_{1}+\lambda\cdot\mathcal{L}_{\text{SSIM}},$
    where $\widehat{\bm{C}}$ is the rendered image, $\bm{C}^{\star}$ the ground truth, and $\lambda\in[0,1]$ balances the two loss terms.
  5. 5.
    Adaptive Densification & Pruning. During training, Gaussians are cloned (under-reconstruction) or split (over-reconstruction) based on view-space gradient magnitude, and low-contribution Gaussians are pruned.
B.3 Involved 3DGS Variants
Although 3DGS features a well-designed pipeline and an efficient differentiable rasterization scheme, its densification process still suffers from several limitations at the detail level. For example, the non-uniform distribution of Gaussians can lead to overlapping primitives in dense regions and under-reconstruction in sparse areas. In addition, the optimization tends to favor large Gaussians that quickly cover the scene to achieve fast convergence. This behavior, however, often results in poor reconstruction of high-frequency details and introduces geometric noise.
Consequently, a growing body of work has focused on refining the densification stage of the original 3DGS framework. In this section, we highlight two such methods—Atom-GS [13] and Mini-Splatting (-D) [12]—which specifically enhance the densification process. While these methods also modify other components such as pruning, we isolate and utilize only the early-stage point cloud generated before 15K training iterations. Applying our compaction method to these two distinct variants further illustrates its generality.
It is important to note that our approach is broadly applicable to any point cloud produced by 3DGS-style pipelines, as long as each Gaussian retains its mean, covariance, and opacity, along with a reasonable optimization strategy for opacity and color.
Atom-GS
To address the tendency of 3DGS to rely on large Gaussians for rapid scene coverage, AtomGS replaces the free-form Gaussian optimization in 3DGS with a structured formulation based on atom Gaussians—uniform, isotropic primitives that better adhere to the underlying scene geometry. In addition, AtomGS introduces geometry-guided optimization to further enhance the spatial fidelity of the reconstructed scene. AtomGS builds upon two key techniques:
  1. 1.
    Atomized Proliferation.
    • Constrains Gaussians in high-detail regions into fixed-size, isotropic Atom Gaussians.
    • Uses a geometric schedule to refine scale and enhance surface coverage over training iterations.
    • Merges Gaussians in flat areas to reduce redundancy.
  2. 2.
    Geometry-Guided Optimization.
    • Introduces Edge-Aware Normal Loss by weighting normal smoothness using an edge map derived from the RGB image.
    • Preserves sharp features while smoothing flat regions, improving alignment with the true 3D geometry.
Mini-Splatting (-D)
The original 3DGS often suffers from non-uniform Gaussian distributions, resulting in excessive overlap in dense regions and insufficient reconstruction in sparse areas. Mini-Splatting addresses this issue by reorganizing the spatial distribution of Gaussians through a two-stage pipeline consisting of densification and simplification. Its goal is to reduce the total number of Gaussians without compromising rendering quality. Mini-Splatting-D is a variant of Mini-Splatting, tailored for quality-oriented rendering. While it follows a similar overall pipeline, the key difference lies in the omission of the simplification stage, allowing the method to focus entirely on maximizing rendering quality. In our implementation, we focus solely on the densification process used in our pipeline.
  • Blur Split: Identifies Gaussians that dominate blurry regions (via screen-space analysis) and splits them to recover fine detail.
  • Depth Reinitialization: Reprojects depth maps derived from midpoints between ray-ellipsoid intersections to generate dense and uniform 3D points for reinitialization.
Beyond evaluating compatibility with different 3DGS densification variants, we compare our method against recent state-of-the-art compaction techniques. Specifically, we consider two representative baselines: LightGaussian [11] and PUP-3DGS [15]. A brief overview of each method is provided below.
LightGaussian
Original 3DGS overpopulates Gaussians during densification, causing storage bloat and rendering inefficiency. LightGaussian prunes redundant Gaussians based on a global significance score (computed via hit count, opacity, and normalized volume), followed by fine-tuning to recover lost details. The steps include:
  1. 1.
    Compute significance for each Gaussian by a score based on the volume of the Gaussian sphere and the number of covered pixels.
  2. 2.
    Prune low-score Gaussians.
  3. 3.
    Fine-tune retained Gaussians with photometric loss.
PUP-3DGS
LightGaussian employs a visibility-aware importance score, which tends to favor pruning small Gaussians while retaining larger background primitives. In contrast, PUP-3DGS introduces a mathematically principled importance measure based on second-order derivatives, allowing it to preserve small but geometrically critical Gaussians in foreground regions. The technical steps are:
  1. 1.
    Fisher Approximation. Computes the gradients of the Gaussian parameters to approximate the Hessian matrix, which is then used to derive the sensitivity score.
  2. 2.
    Iterative Pruning. Rank Gaussians by sensitivity scores, prune low-sensitivity ones, and fine-tune iteratively to recover lost fidelity.
Appendix C Sensitivity Analysis
C.1 Choice of KD-tree Depth
In our algorithm, the KD-tree depth $d$ (or equivalently, the number $s$ of Gaussians per KD block) is treated as a hyperparameter. We investigate how rendering quality varies with different values of $d$ on the Truck scene. The results are shown in Figure 6.
Figure 6: Sensitivity Analysis of $d$. Our GHAP algorithm exhibits stable rendering performance across varying KD-tree depths $d$.
We observe that, within a reasonable range, the KD-tree depth $d$ has minimal impact on rendering quality. In our experiments on the Truck scene, setting $d<6$ results in distance matrices that are too large to fit into GPU memory, causing memory overflow. Conversely, when $d>16$, the number of partitions becomes excessively large, and most blocks contain one or zero Gaussians, rendering the algorithm inapplicable.
C.2 Sampling Iteration
Since most 3DGS-based algorithms complete the densification stage before 15K training iterations, we apply our sampling procedure after this point to achieve a target retention ratio. To study the effect of sampling at different post-densification stages, we evaluate our method at various iterations beyond 15K. The experimental results are presented in Figure 7.
Figure 7: Sensitivity Analysis of Sampling Iteration. The rendering performance of our GHAP algorithm has negligible loss by variations in sampling iteration.
As shown in the results, sampling at any iteration beyond 15K has minimal impact on the final rendering quality. These results indicate that continuing optimization beyond 15K provides diminishing returns. Moreover, once sampling is performed, only a small number of fine-tuning iterations are needed for the compacted model to converge. Based on these observations, we recommend applying sampling uniformly at the 15K iteration, followed by an additional 15K fine-tuning iterations for all experiments.
Appendix D Additional Scene Visualizations
We report a more comprehensive set of results for the plug-in experiments in Table 4, including two additional metrics: the number of Gaussians (in thousands) and rendering speed measured in FPS. As shown in the updated results, our compaction method significantly improves FPS compared to the original models, often yielding over a 2$\times$ speedup in rendering performance.
In addition to the previously shown Figure 4, we present more detailed qualitative comparisons across multiple scenes in Figure 8. Our method consistently preserves the visual quality of the original models. Notably, when applied to stronger 3DGS variants with improved densification strategies—such as Mini-Splatting-D—our compaction framework performs even better. This is reflected in the fact that, after compaction, Mini-Splatting-D often achieves higher rendering quality than the original 3DGS baseline.
Table 4: Additional information in plug-in experiment.
Method-Compact Rate $\rho$ Tank&Temples MipNeRF-360 Deep Blendering
SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS SSIM PSNR LPIPS $k$ Gaussians FPS
3DGS-30k 10% 0.818 23.312 0.242 157 341 0.764 26.404 0.314 272 272 0.905 29.647 0.264 248 290
20% 0.835 23.615 0.212 313 300 0.788 26.973 0.275 545 230 0.907 29.864 0.252 496 237
100% 0.853 23.785 0.169 1577 168 0.813 27.554 0.221 2625 126 0.907 29.816 0.238 2476 126
Mini-Splatting-D 10% 0.835 23.232 0.198 147 564 0.802 27.090 0.25 318 389 0.909 30.042 0.254 156 549
20% 0.855 23.403 0.171 293 526 0.821 27.310 0.214 636 299 0.912 30.17 0.238 318 489
100% 0.848 23.338 0.14 1473 209 0.832 27.486 0.176 3177 111 0.907 29.980 0.211 1597 197
Atom-GS 10% 0.793 22.988 0.274 426 401 0.764 26.535 0.307 468 304 0.899 29.347 0.282 464 419
20% 0.812 23.282 0.24 852 283 0.788 27.025 0.269 937 236 0.902 29.347 0.268 925 323
100% 0.814 23.289 0.235 4274 87 0.796 27.135 0.251 4699 86 0.902 29.314 0.267 4625 121
Figure 8: More scenes visualization. Visual comparisons across additional scenes before and after compaction.